# date20082024

### v1

@@ Abstract
In cavity quantum electrodynamics (QED), emitters and a resonator are coupled together to enable precise studies of quantum light-matter interactions. Over the past few decades, this has led to a variety of quantum technologies such as more precise inertial sensors, clocks, memories, controllable qubits, and quantum simulators [1, 2, 3, 4, 5, 6, 7]. Furthermore, the intrinsically dissipative nature of cavity QED platforms makes them a natural testbed for exploring driven-dissipative phenomena in open quantum systems as well as equilibrium and non-equilibrium phase transitions in quantum optics [8, 9, 10, 11, 12]. One such model, the so-called cooperative resonance fluorescence (CRF) model, concerns the behavior of coherently driven emitters in the presence of collective dissipation (superradiance). Despite tremendous interest [13, 14, 15, 16, 17, 18], this model has yet to be realized in a clean experimental system. Here we provide an observation of the continuous superradiant phase transition predicted in the CRF model using an ensemble of ultracold 88Sr atoms coupled to a driven high-finesse optical cavity on a long-lived optical transition. Below a critical drive, atoms quickly reach a steady state determined by the self-balancing of the drive and the collective dissipation. The steady state possesses a macroscopic dipole moment and corresponds to a superradiant phase. Above a critical drive strength, the atoms undergo persistent Rabi-like oscillations until other decoherence processes kick in. In fact, our platform also allows us to witness the change of this phase transition from second to first order induced by single-particle spontaneous emission, which pushes the system towards a different steady state. Our observations are a first step towards finer control of driven-dissipative systems, which have been predicted to generate quantum states that can be harnessed for quantum information processing and in particular quantum sensing [19, 2, 20, 21, 22].
@@
Fig. 1: Experimental schematics and the phase diagram of the cooperative resonance fluorescence (CRF) model. a, ${}^{88}\mathrm{Sr}$ atoms (green) are trapped in an optical lattice (omitted in the figure for simplicity) supported by a high-finesse optical cavity (grey mirrors) after laser cooling. Atoms couple to the cavity with a r.m.s. single-photon Rabi frequency $2g$. They have a spontaneous decay rate $\gamma$. The cavity has a FWHM linewidth of $\kappa$. We inject 689689689~{}689nm light into the cavity, which would establish a drive field (red) with Rabi frequency $\Omega_{d}$ inside the cavity. In response, atoms emit a superradiant field (blue) at the same wavelength, characterized by Rabi frequency $\Omega_{\mathrm{SR}}=i\tfrac{4g^{2}}{\kappa}\langle{\hat{J}}^{-}\rangle$. The coherent addition of the two fields (purple) leaks through the cavity mirror at a rate $\kappa$ and is monitored on a detector (brown). b, Energy level diagram of experimentally relevant frequency scales. The optical cavity resonance frequency $\omega_{c}$ and the atomic drive frequency $\omega_{d}$ are both resonant with the ${}^{1}\textrm{S}_{0}-^{3}\!\text{P}_{1}(m_{J}=0)$ transition. c, Phase diagram of the CRF Model. The upper two panels show cuts of the Bloch sphere illustration in the y-z plane for the superradiant phase (left) and the normal phase (right), with the dashed lines indicating the representative points on the phase diagram on the lower row. The external drive $\Omega_{d}$ (red) acts like a torque that tries to rotate the collective Bloch vector $\vec{J}$ (green) counter-clockwise; while the superradiant field from the atoms $\Omega_{\mathrm{SR}}$ (blue) acts like a torque, with strength proportional to the collective atomic dipole moment $J^{-}$, that always tries to bring the Bloch vector back to the ground state (south pole). In the superradiant phase ($\Omega_{d}<\Omega_{c}^{h}=NC\gamma/2=2Ng^{2}/\kappa$), the drive and the dissipation in the form of superradiance balance each other and atoms in the steady state reach a fixed spin-projection below the equator, $J_{z}<0$. Above the critical point, $\Omega_{d}>\Omega_{c}^{h}$, normal phase, the superradiant field from the atoms is no longer strong enough to cancel the drive field, so atoms start to Rabi flop and their spin-projection $J_{z}$ time-averages to zero. The bottom row shows the long time steady state’s normalized spin projection $J_{z}/(N/2)$ as a function of normalized Rabi frequency $\Omega_{d}/\Omega_{c}^{h}$. $J_{z}$ can be identified as an indicator of a second-order phase transition. Spontaneous emission modifies the steady state behaviour of the transition and changes it to a first order transition at a smaller critical drive (orange curve).
@@ Introduction
Here we report an observation of the second-order superradiant transition in CRF by resonantly driving an ensemble of 88Sr atoms on a narrow linewidth transition inside of a resonant high finesse cavity. We identify and distinguish the two predicted phases by observing both spin and light degrees of freedom as the drive strength is varied. Central to our observations are two key elements of the experiment. First, the use of a high-finesse optical cavity allows us to operate with a spatially dilute ensemble of atoms with negligible dipolar or contact interactions; the cavity also provides well-defined phase matching to a single optical mode, ensuring the atoms are always in a fully collective regime. Second, the interrogation of a narrow-linewidth atomic transition enables us to observe the collective physics of CRF on an experimentally accessible timescale. The clean separation of timescales between the collective decay and single-particle spontaneous decay also allows us to observe how spontaneous emission slowly steers the system towards a qualitatively different steady state that displays a discontinuous transition in both spin and light observables. This physics is related to prior studies in optical bistability [54, 55, 56, 57, 58], although past experimental works have focused mostly on the properties of the light, such as hysteretic transmission, and on time scales where spontaneous emission is important. Our work paves the way for experimental verification of new symmetries in open quantum systems [59, 60, 61, 21] and opens the door for exploring new ways of generating spin-squeezing [21] and spectroscopy on ultranarrow clock transitions [62, 63].
@@
Here, $\hat{\rho}$ is the density matrix of the full atom-cavity system, $\hat{J}^{-}=\sum_{i=1}^{N}\hat{s}^{-}_{i}$ and $\hat{J}^{+}=(\hat{J}^{-})^{\dagger}$ are, respectively, collective spin lowering and raising operators that quantify the atomic coherence between the two-level systems, the operator $\hat{a}^{\dagger}$ ($\hat{a}$) creates (destroys) one photon in the cavity mode, and the Lindbladian super-operator $\mathcal{L}_{c}(\hat{\rho})=\hat{a}\hat{\rho}\hat{a}^{\dagger}-\{\hat{a}^{\dagger}\hat{a},\hat{\rho}\}/2$ describes photon emission through the cavity. The first term in the Hamiltonian $\hat{H}$ is the Tavis-Cummings atom-cavity interaction, and the second term represents the laser drive with a strength quantified by the Rabi frequency $\Omega_{d}=\alpha\times 2g$, where $\alpha$ is the coherent state amplitude that would be established inside the cavity by the drive if there were no atoms. For now, we have omitted single-particle spontaneous emission, which happens on a slower timescale set by $\gamma^{-1}$.
Fig. 2: Observing the second-order superradiant phase transition. a,b Steady-state spin projection $J_{z}$ after $9.3~{}\mu$s of the drive as a function of Rabi frequency $\Omega_{d}$ (a) and normalized Rabi frequency $\Omega_{d}/\Omega_{c}$ with $\Omega_{c}=0.35NC\gamma$ (b). Regions of gray here and hereafter indicate that our measurement of spin-projection $J_{z}$ does not accurately capture the real $J_{z}$ value of our system at the end of the drive (see main text). The three differently colored traces are for three different total atom numbers $N$ with the error bar denoting the standard deviation. In b, below the critical drive, the three data traces at different $N$ collapse on top of each other after normalizing to the collective decay rate $\Omega_{c}$, which at the same time align well with the theory prediction (solid black curve) that $J_{z}=0$ at $\Omega_{d}/\Omega_{c}=1$. This certifies the observation of the superradiant phase. c, Transmission of the cavity with atoms inside normalized to that of a cavity without atoms $T$ as a function of normalized Rabi frequency $\Omega_{d}/\Omega_{c}$. The transmitted power is averaged over the last $5~{}\mu$s of the drive duration. The measured transmission is consistent with 0 below the critical Rabi drive $\Omega_{c}$, which indicates the complete cancellation of the external drive field by the superradiantly emitted field; the fact that transmission becomes finite above the critical drive $\Omega_{c}$ confirms that the superradiant field is no longer large enough to cancel the drive in the normal phase. Error bars in the plots, and hereafter, represent standard error of the mean unless otherwise noted.
@@
The properties of the steady state of the system, $d\hat{\rho}_{ss}/dt=0$, can be understood by examining the mean-field equation of motion for the intracavity field $a=\langle\hat{a}\rangle$ and the atomic coherence $J^{-}=\langle\hat{J}^{-}\rangle$:
@@
where $J_{z}=\langle\hat{J}_{z}\rangle=\sum_{i}\langle\hat{s}^{z}_{i}\rangle$ represents the collective spin projection. In the equation for the field, the first term on the right hand side describes the external drive, the second term describes the superradiant field radiated into the cavity by the collective atomic dipole moment with strength proportional to $J^{-}$, and the last term accounts for the decay of field from the cavity. The other two equations describe how the intracavity field rotates the collective Bloch vector.
@@
Fig. 3: Modification of critical behaviour by single-particle spontaneous emission. a, Bloch sphere illustration of the superradiant phase steady state without spontaneous emission (left) and with spontaneous emission (right). On the left, the steady state Bloch vector $\vec{J}$ stays in the collective manifold where the torque coming from the external drive $\Omega_{d}$ (red arrow) and superradiant field $\Omega_{\textrm{SR}}$ (blue arrow) can be balanced. On the right, the steady state Bloch vector $\vec{J}$ is no longer in the collective manifold. The Bloch vector is rotated up from the south pole to increase the coherence $J^{-}$ such that the superradiant field can still cancel the external drive. As a result, the steady state $J_{z}$ on the right is always higher than that on the left. b, Spin projection $J_{z}$ as a function of drive duration $t$. Traces with different colors represent different drive strengths with fixed atom number. The solid lines represent the results of numerical simulation. The tick marks on the top horizontal axis indicate the drive duration as a function of multiples of $\tau=\gamma^{-1}=21~{}\mu$s, the 1/$e$ lifetime of the excited state. Traces with $\Omega_{d}/\Omega_{c}=0.50,0.82$ have $J_{z}$ slowly relax above 0 on a time scale of order the excited state lifetime, while the traces with $\Omega_{d}/\Omega_{c}=0.19,0.31$ have $J_{z}$ remain below 0 and stay in the superradiant phase. The vertical colored dashed lines indicate the times at which corresponding colored data traces are presented in c, d. c, Spin projection $J_{z}$ as a function of $\Omega_{d}/\Omega_{c}$ with fixed atom number. Different colors indicate different drive durations $t$ before the measurement of $J_{z}$. The critical point, where $J_{z}$ crosses 0, reduces to smaller drive values as drive duration increases, and the continuous second order phase transition ($t=9~{}\mu$s green curve) turns into a first order transition ($t=300~{}\mu$s, red curve), where a clear jump is observed. d, Fractional transmission $T$ as a function of $\Omega_{d}/\Omega_{c}$. The drive strength at which transmission becomes non-zero gradually shifts to a lower value as the drive duration increases, which aligns with the critical point suggested by the $J_{z}$ measurement. The green curve ($t=9~{}\mu$s) displays a smooth transition while the red curve with a longer drive duration ($t=300~{}\mu$s) displays a discontinuity at the critical point.
@@
We show the (normalized) spin projection $J_{z}/(N/2)$ as a function of Rabi frequency $\Omega_{d}$ for three different values of the atom number in Fig. 2a. After normalizing the Rabi frequency by $\Omega_{c}=0.70\Omega_{c}^{h}=0.35NC\gamma$, we find that the curves in the region $\Omega_{d}/\Omega_{c}<1$ collapse on top of each other, certifying the collective nature of the steady state. The numerical factor 0.70 accounts mainly for inhomogenous atom-cavity couplings, which reduce the critical Rabi drive but do not change the second-order nature of the transition (see Methods, SI). The collapsed data is in good agreement with numerical simulations (solid black curve), consistent with the prediction of the superradiant phase. In the region where $\Omega_{d}>\Omega_{c}$, our shelving procedure does not properly capture the spin projection $J_{z}$ at the end of the hold period. This is because in this regime, the cavity is populated with a macroscopic field that rapidly transfers atoms from $\ket{\downarrow}$ to $\ket{\uparrow}$ where they are then subsequently shelved, leading to a miss-inferred value of $J_{z}>0$. To emphasize this, we shade the regions with $J_{z}>0$ in grey in Figures 2 to 5. Note, however, that for smaller atom number, we expect the intracavity field to be relatively small just above the transition, allowing the shelving method to provide a more accurate inference of $J_{z}$. Indeed, we observe in Fig. 2b that for $N=3.4\times 10^{3}$, the measured value of $J_{z}$ exhibits an abrupt change of behavior at $\Omega_{d}=\Omega_{c}$.
@@
By driving the system for longer times, we can explore the effect of a different type of dissipation: spontaneous emission, which is a non-collective effect described by another Lindblad term added to the master equation Eq.(1), $\gamma\mathcal{L}_{se}(\hat{\rho})=\gamma\sum_{i}(\hat{s}_{i}^{-}\hat{\rho}\hat{s}_{i}^{+}-\{\hat{s}_{i}^{+}\hat{s}_{i}^{-},\hat{\rho}\}/2)$. Unlike superradiant decay, which preserves the total length of the Bloch vector $J\approx\sqrt{|J^{-}|^{2}+J_{z}^{2}}$, spontaneous emission shortens $J$. If the system starts in the steady state of the superradiant phase in the CRF model, the atomic coherence $J^{-}$ is shortened by spontaneous emission and in turn radiates a smaller field. This leads to a non-zero net field inside the cavity due to the imbalance of the drive and superradiant field. The Bloch vector then rotates upwards, which restores $J^{-}$ to its previous value but introduces an effective upwards force on $J_{z}$ that scales with the drive strength $\Omega_{d}$ (see SI). At the same time, the decay to the ground state induced by spontaneous emission leads to a competing process that pushes $J_{z}$ down towards $-N/2$. This competition leads to two different behaviours depending on the drive strength. For strong drive strengths, the upwards force causes $J_{z}$ to slowly relax towards 0 while holding $J^{-}$ constant. Once $J_{z}$ reaches 0, the atoms start to undergo Rabi-like oscillation, and the cavity gets quickly populated by photons, indicative of the normal phase despite having started in the CRF superradiant phase. However, when $\Omega_{d}$ is small enough, since the upwards force on $J_{z}$ scales with $\Omega_{d}$, $J_{z}$ will not relax all the way towards 0. Instead, the Bloch vector arrives at a new steady state where the upwards and downwards forces balance, and the superradiant and drive fields retain a near-perfect cancellation despite the reduction in spin length $J$ (see SI). This corresponds to the superradiant phase (see Fig 3a), which is analogous to the “cooperative branch” in the bistability literature [69, 54]. These two regimes are separated by a new critical point $\Omega_{c}^{\prime}$, and the spin projection $J_{z}$ will now display a discontinuous jump at this point as opposed to the continuous transition in the CRF model [17, 70] (also see Fig.1c).
@@
We show traces of the spin projection $J_{z}$ as a function of drive duration $t$ for different drive strengths in Fig. 3b. At low drives ($\Omega_{d}/\Omega_{c}=0.19,0.31)$, $J_{z}$ reaches its fully collective superradiant steady state at short times, and at longer times the Bloch vector still remains below the south pole with a modified value. As the drive strength is increased ($\Omega_{d}/\Omega_{c}=0.50,0.82$), the spin projection $J_{z}$ remains below the equator on timescales much smaller than the timescale of the excited state lifetime $\tau=\gamma^{-1}=21~{}\mu$s. For longer times, however, it slowly moves towards 00, consistent with theoretical expectations. At even larger drives ($\Omega_{d}/\Omega_{c}=1.4$) the inversion quickly rises above 0, suggesting the systems are in the normal phase both at short and long times.
@@
We characterize the gradual change in $J_{z}$ on the timescale of $\tau$ by providing snapshots of $J_{z}$ as a function of normalized Rabi frequency $\Omega_{d}/\Omega_{c}$ for different drive durations in Fig. 3c. At short times ($9~{}\mu$s), we observe a continuous superradiant phase transition, consistent with Fig. 2b. At longer times ($t\gtrsim\tau$), spontaneous emission becomes important and noticeably shifts the transition point, characterized by the point where $J_{z}$ crosses 0, to a lower value of $\Omega_{d}/\Omega_{c}$. By $t=100~{}\mu$s, the system has already equilibrated to a new steady state, and the spin projection clearly displays a discontinuous jump at $\Omega_{d}/\Omega_{c}=0.4$. These observations can be correlated with the results of fractional transmission, $T$, through the cavity, shown in Fig. 3d as a function of $\Omega_{d}/\Omega_{c}$ for different drive durations. Again, the transition points where the transmission rises above 0 coincide with the points at which $J_{z}$ reaches 0 in Fig. 3c.

### v2

# date26082024

@@ Abstract
In cavity quantum electrodynamics (QED), emitters and a resonator are coupled together to enable precise studies of quantum light-matter interactions. Over the past few decades, this has led to a variety of quantum technologies such as more precise inertial sensors, clocks, memories, controllable qubits, and quantum simulators [1, 2, 3, 4, 5, 6, 7]. Furthermore, the intrinsically dissipative nature of cavity-QED platforms makes them a natural testbed for exploring driven-dissipative phenomena in open quantum systems as well as equilibrium and non-equilibrium phase transitions in quantum optics [8, 9, 10, 11, 12]. One such model, the so-called cooperative resonance fluorescence (CRF) model, concerns the behavior of coherently driven emitters in the presence of collective dissipation (superradiance). Despite tremendous interest [13, 14, 15, 16, 17, 18], this model has yet to be realized in a clean experimental system. Here we provide an observation of the continuous superradiant phase transition predicted in the CRF model using an ensemble of ultracold 88Sr atoms coupled to a driven high-finesse optical cavity on a long-lived optical transition. Below a critical drive, atoms quickly reach a steady state determined by the self-balancing of the drive and the collective dissipation. The steady state possesses a macroscopic dipole moment and corresponds to a superradiant phase. Above a critical drive strength, the atoms undergo persistent Rabi-like oscillations until other decoherence processes kick in. In fact, our platform also allows us to witness the change of this phase transition from second to first order induced by single-particle spontaneous emission, which pushes the system towards a different steady state. Our observations are a first step towards finer control of driven-dissipative systems, which have been predicted to generate quantum states that can be harnessed for quantum information processing and in particular quantum sensing [19, 2, 20, 21, 22].
@@
Fig. 1: Experimental schematics and the phase diagram of the CRF model. a, ${}^{88}\mathrm{Sr}$ atoms (green) are trapped in an optical lattice (omitted in the figure for simplicity) supported by a high-finesse optical cavity (grey mirrors) after laser cooling. The atoms couple to the cavity with a r.m.s. single-photon Rabi frequency $2g$. The excited state has a spontaneous decay rate $\gamma$. The cavity has a FWHM linewidth of $\kappa$. We inject 689689689~{}689nm light into the cavity, which would establish a drive field (red) with Rabi frequency $\Omega_{d}$ inside the cavity. In response, the atoms emit a superradiant field (blue) at the same wavelength, characterized by Rabi frequency $\Omega_{\mathrm{SR}}=i\tfrac{4g^{2}}{\kappa}\langle{\hat{J}}^{-}\rangle$. The coherent addition of the two fields (purple) leaks through the cavity mirror at a rate $\kappa$ and is monitored on a detector (brown). b, Energy level diagram of experimentally relevant frequency scales. The optical cavity resonance frequency $\omega_{c}$ and the atomic drive frequency $\omega_{d}$ are both resonant with the ${}^{1}\textrm{S}_{0}-^{3}\!\text{P}_{1}(m_{J}=0)$ transition. c, Phase diagram of the CRF Model. The upper two panels show cuts of the Bloch sphere illustration in the y-z plane for the superradiant phase (left) and the normal phase (right), with the dashed lines indicating the representative points on the phase diagram on the lower row. The external drive $\Omega_{d}$ (red) acts like a torque that tries to rotate the collective Bloch vector $\vec{J}$ (green) counter-clockwise. The superradiant field from the atoms $\Omega_{\mathrm{SR}}$ (blue) acts like a torque, with strength proportional to the collective atomic dipole moment $J^{-}$, that always tries to bring the Bloch vector back to the ground state (south pole). In the superradiant phase ($\Omega_{d}<\Omega_{c}^{h}=NC\gamma/2=2Ng^{2}/\kappa$), the drive and the dissipation in the form of superradiance balance each other, and the atoms reach a steady state with a fixed spin-projection below the equator, $J_{z}<0$. Above the critical point ($\Omega_{d}>\Omega_{c}^{h}$), i.e., the normal phase, the superradiant field from the atoms is no longer strong enough to cancel the drive field, so the Bloch vector starts to Rabi flop, and the spin-projection $J_{z}$ time-averages to zero. The bottom row shows the steady state’s normalized spin projection $J_{z}/(N/2)$ as a function of normalized Rabi frequency $\Omega_{d}/\Omega_{c}^{h}$. $J_{z}$ can be identified as an indicator of a second-order phase transition. Spontaneous emission modifies the steady state behaviour of the transition and changes it to a first order transition at a smaller critical drive (orange curve).
@@ Introduction
Here we report an observation of the second-order superradiant transition in CRF by resonantly driving an ensemble of 88Sr atoms on a narrow linewidth transition inside of a resonant high-finesse cavity. We identify and distinguish the two predicted phases by observing both spin and light degrees of freedom as the drive strength is varied. Central to our observations are two key elements of the experiment. First, the use of a high-finesse optical cavity allows us to operate with a spatially dilute ensemble of atoms with negligible dipolar or contact interactions; the cavity also provides well-defined phase matching to a single optical mode, ensuring the atoms are always in a fully collective regime. Second, the interrogation of a narrow-linewidth atomic transition enables us to observe the collective physics of CRF on an experimentally accessible timescale. The clean separation of timescales between the collective decay and single-particle spontaneous decay also allows us to observe how spontaneous emission slowly steers the system towards a qualitatively different steady state that displays a discontinuous transition in both spin and light observables. This physics is related to prior studies in optical bistability [54, 55, 56, 57, 58], although past experimental works have focused mostly on the properties of the light, such as hysteretic transmission, and on timescales where spontaneous emission is important. Our work paves the way for experimental verification of new symmetries in open quantum systems [59, 60, 61, 21] and opens the door for exploring new ways of generating spin-squeezing [21] and spectroscopy on ultranarrow clock transitions [62, 63].
@@
Here, $\hat{\rho}$ is the density matrix of the full atom-cavity system, $\hat{J}^{-}=\sum_{i=1}^{N}\hat{s}^{-}_{i}$ and $\hat{J}^{+}=(\hat{J}^{-})^{\dagger}$ are, respectively, collective spin lowering and raising operators that quantify the collective atomic coherence along the two-level system, the operator $\hat{a}^{\dagger}$ ($\hat{a}$) creates (destroys) one photon in the cavity mode, and the Lindbladian superoperator $\mathcal{L}_{c}(\hat{\rho})=\hat{a}\hat{\rho}\hat{a}^{\dagger}-\{\hat{a}^{\dagger}\hat{a},\hat{\rho}\}/2$ describes photon emission through the cavity. The first term in the Hamiltonian $\hat{H}$ is the Tavis-Cummings atom-cavity interaction, and the second term represents the laser drive with a strength quantified by the Rabi frequency $\Omega_{d}=\alpha\times 2g$, where $\alpha$ is the coherent state amplitude that would be established inside the cavity by the drive without atoms. For now, we have omitted single-particle spontaneous emission, which happens on a slower timescale set by $\gamma^{-1}$.
Fig. 2: Observing the second-order superradiant phase transition. a,b Steady-state spin projection $J_{z}$ after a $9.3~{}\mu$s drive as a function of (a) Rabi frequency $\Omega_{d}$ and (b) normalized Rabi frequency $\Omega_{d}/\Omega_{c}$ with $\Omega_{c}=0.35NC\gamma$. Regions of gray hereafter indicate that our measurement of spin-projection $J_{z}$ does not accurately capture the real $J_{z}$ value of our system at the end of the drive duration (see main text). The three differently colored datasets are for three different total atom numbers $N$ with the error bars denoting the standard deviation. In b, below the critical drive, the three sets at different $N$ collapse on top of each other after normalizing to the collective decay rate $\Omega_{c}$, which at the same time align well with the theory prediction (solid black curve). c, The transmission $T$ of the cavity with atoms inside normalized to that of a cavity without atoms as a function of normalized Rabi frequency $\Omega_{d}/\Omega_{c}$. The transmitted power is averaged over the last $5~{}\mu$s of the drive duration. The measured transmission is consistent with 0 below the critical Rabi drive $\Omega_{c}$, which indicates the complete cancellation of the external drive field by the superradiantly emitted field; the fact that transmission becomes finite above the critical drive $\Omega_{c}$ confirms that the superradiant field is no longer large enough to cancel the drive in the normal phase. Error bars in the plots, and hereafter, represent the standard error of the mean unless otherwise noted.
@@
The properties of the steady state of the system, which satisfies $d\hat{\rho}_{ss}/dt=0$, can be understood by examining the mean-field equation of motion for the intracavity field $a=\langle\hat{a}\rangle$ and the atomic coherence $J^{-}=\langle\hat{J}^{-}\rangle$:
@@
where $J_{z}=\langle\hat{J}_{z}\rangle=\sum_{i}\langle\hat{s}^{z}_{i}\rangle$ represents the collective spin projection. In the equation for the field, the first term on the right hand side describes the external drive, the second term describes the superradiant field radiated into the cavity by the collective atomic dipole moment with strength proportional to $J^{-}$, and the last term accounts for the cavity field decay. The other two equations describe how the intracavity field rotates the collective Bloch vector.
@@
Fig. 3: Modification of critical behaviour by single-particle spontaneous emission. a, Bloch sphere illustration of the superradiant phase steady state without (left) and with (right) spontaneous emission. On the left, the steady state Bloch vector $\vec{J}$ stays in the collective manifold where the torque coming from the external drive $\Omega_{d}$ (red arrow) and superradiant field $\Omega_{\textrm{SR}}$ (blue arrow) can be balanced. On the right, the steady state Bloch vector $\vec{J}$ is no longer in the collective manifold. The Bloch vector is rotated up from the south pole to increase the coherence $J^{-}$ such that the superradiant field can still cancel the external drive. As a result, the steady state $J_{z}$ on the right is always higher than that on the left. b, The spin projection $J_{z}$ as a function of drive duration $t$. Traces with different colors represent different drive strengths with fixed atom number. The solid lines represent the results of numerical simulation. The tick marks on the top horizontal axis indicate the drive duration as a function of multiples of $\tau=\gamma^{-1}=21~{}\mu$s, the 1/$e$ lifetime of the excited state. Traces with $\Omega_{d}/\Omega_{c}=0.50,0.82$ have $J_{z}$ slowly relax above 0 on a timescale set by the excited state lifetime, while the traces with $\Omega_{d}/\Omega_{c}=0.19,0.31$ have $J_{z}$ remain below 0 and stay in the superradiant phase. The coloured vertical dashed lines indicate the drive durations corresponding to the data with matching colours in c, d. c, $J_{z}$ as a function of $\Omega_{d}/\Omega_{c}$ with fixed atom number. Different colors indicate different drive durations $t$ before the measurement of $J_{z}$. The critical point, where $J_{z}$ crosses 0, reduces to smaller drive values as drive duration increases, and the continuous second order phase transition ($t=9~{}\mu$s, green curve) turns into a first order transition ($t=300~{}\mu$s, red curve), where a clear jump is observed. d, The fractional transmission $T$ as a function of $\Omega_{d}/\Omega_{c}$. The drive strength at which transmission becomes non-zero gradually shifts to a lower value as the drive duration increases, which aligns with the critical point suggested by the $J_{z}$ measurement. The green curve ($t=9~{}\mu$s) displays a smooth transition while the red curve with a longer drive duration ($t=300~{}\mu$s) displays a discontinuity at the critical point.
@@
We show the (normalized) spin projection $J_{z}/(N/2)$ as a function of Rabi frequency $\Omega_{d}$ for three different atom numbers in Fig. 2a. After normalizing the Rabi frequency by $\Omega_{c}=0.70\Omega_{c}^{h}=0.35NC\gamma$, we find that the curves in the region $\Omega_{d}/\Omega_{c}<1$ collapse on top of each other, certifying the collective nature of the steady state. The numerical factor 0.70 mainly accounts for inhomogenous atom-cavity coupling, which reduces the critical Rabi drive but does not change the second-order nature of the transition (see Methods, SI). The collapsed data is in good agreement with numerical simulations (solid black curve), consistent with the prediction of the superradiant phase. In the region where $\Omega_{d}>\Omega_{c}$, our shelving procedure does not properly capture the spin projection $J_{z}$ at the end of the hold period. This is because in this regime, the cavity is populated with a macroscopic field that rapidly transfers atoms from $\ket{\downarrow}$ to $\ket{\uparrow}$ where they are then subsequently shelved, leading to a mis-inferred value of $J_{z}>0$. To emphasize this, we shade the regions with $J_{z}>0$ in grey in Figures 2 to 5. Note, however, that for smaller atom number, we expect the intracavity field to be relatively small just above the transition, allowing the shelving method to provide a more accurate inference of $J_{z}$. Indeed, we observe in Fig. 2b that for $N=3.4\times 10^{3}$, the measured value of $J_{z}$ exhibits an abrupt change of behavior at $\Omega_{d}=\Omega_{c}$.
@@
By driving the system for longer times, we can explore the effect of a different type of dissipation: spontaneous emission, which is a non-collective effect described by another Lindblad term added to the master equation Eq. (1), $\gamma\mathcal{L}_{se}(\hat{\rho})=\gamma\sum_{i}(\hat{s}_{i}^{-}\hat{\rho}\hat{s}_{i}^{+}-\{\hat{s}_{i}^{+}\hat{s}_{i}^{-},\hat{\rho}\}/2)$. Unlike superradiant decay, which preserves the total length of the Bloch vector $J\approx\sqrt{|J^{-}|^{2}+J_{z}^{2}}$, spontaneous emission shortens $J$. If the system starts in the steady state of the superradiant phase in the CRF model, the atomic coherence $J^{-}$ is shortened by spontaneous emission and in turn radiates a smaller field. This leads to a non-zero net field inside the cavity due to the imbalance of the drive and superradiant field. The Bloch vector then rotates upwards, which restores $J^{-}$ to its previous value but introduces an effective upwards force on $J_{z}$ that scales with the drive strength $\Omega_{d}$ (see SI). At the same time, the decay to the ground state induced by spontaneous emission leads to a competing process that pushes $J_{z}$ down towards $-N/2$. This competition leads to two different behaviours depending on the drive strength. For strong drive strengths, the upwards force causes $J_{z}$ to slowly relax towards 0 while holding $J^{-}$ constant. Once $J_{z}$ reaches 0, the atoms start to undergo Rabi-like oscillation, and the cavity gets quickly populated by photons, indicative of the normal phase despite having started in the CRF superradiant phase. However, when $\Omega_{d}$ is small enough, since the upwards force on $J_{z}$ scales with $\Omega_{d}$, $J_{z}$ will not relax all the way towards 0. Instead, the Bloch vector arrives at a new steady state where the upwards and downwards forces balance, and the superradiant and drive fields retain a near-perfect cancellation despite the reduction in spin length $J$ (see SI). This corresponds to the superradiant phase (see Fig. 3a), which is analogous to the “cooperative branch” in the bistability literature [69, 54]. These two regimes are separated by a new critical point $\Omega_{c}^{\prime}$, and the spin projection $J_{z}$ will now display a discontinuous jump at this point as opposed to the continuous transition in the CRF model [17, 70] (also see Fig. 1c).
@@
We show traces of the spin projection $J_{z}$ as a function of the drive duration $t$ for different drive strengths in Fig. 3b. At low drives ($\Omega_{d}/\Omega_{c}=0.19,0.31)$, $J_{z}$ reaches its fully collective superradiant steady state at short times, and at longer times the Bloch vector still remains below the south pole with a modified value. As the drive strength is increased ($\Omega_{d}/\Omega_{c}=0.50,0.82$), the spin projection $J_{z}$ remains below the equator on a timescale shorter than the excited state lifetime $\tau=\gamma^{-1}=21~{}\mu$s. For longer times, however, it slowly moves towards 00, consistent with theoretical expectations. At even larger drives ($\Omega_{d}/\Omega_{c}=1.4$) the inversion quickly rises above 0, suggesting the systems are in the normal phase both at short and long times.
@@
We characterize the gradual change in $J_{z}$ over time by providing snapshots of $J_{z}$ as a function of the Rabi frequency $\Omega_{d}$ for different drive durations in Fig. 3c. At short times ($9~{}\mu$s), we observe a continuous superradiant phase transition, consistent with Fig. 2b. At longer times ($t\gtrsim\tau$), spontaneous emission becomes important and noticeably shifts the transition point, characterized by the point where $J_{z}$ crosses 0, to a lower value of $\Omega_{d}$. By $t=100~{}\mu$s, the system has already equilibrated to a new steady state, and the spin projection clearly displays a discontinuous jump at $\Omega_{d}/\Omega_{c}=0.4$. These observations can be correlated with the results of fractional transmission, $T$, through the cavity, shown in Fig. 3d as a function of $\Omega_{d}$ for different drive durations. Again, the transition points where the transmission rises above 0 coincide with the points at which $J_{z}$ reaches 0 in Fig. 3c.